1.     Determine the wavelength of a radio station transmitting a broadcast at 520 Megahertz.

            a)         0.57 m

            b)         5.8 x 105 m

            c)         1.6 x 1011 m

            d)         1.6 x 105 m
Answer #1

520 MHz = 5.2 * 108 Hz

We have a relationship between Wavelength and frequency :

c = f*L

c - velocity of light

f - frequency

L - wavelength

Therefore,

3 * 108 =5.2 * 108 * L

=> L = 3/5.2 = 0.57m

That is the wavelength of the light emitted is 0.57m.